OpenStudy (anonymous):

WAT IS -21/50 DIVIDED BY 9/15 PLEAZE SHOW WORK

OpenStudy (anonymous):

-21/50 DIVIDED BY 9/15 =(-21/50) /( 9/15) in division, just invert the denominator =(-21/50) ( 15/9) =(-21*15)(50*9) reducing them to lowest term = -7/10 ans..